OpenStudy (anonymous):

Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1 Set #2 (0, 4) and (-2, 8) (3, 7) and (-1, -5) The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (binary3i):

slope = ( Y2-Y1)/(X2-X1) now you can do it.

OpenStudy (anonymous):

thats the part i dont know

OpenStudy (binary3i):

point 1(0,4), point 2(-2,8) so X1=0 and Y1=4 X2=-2 and Y2=8 put it in the equation of slope and slope becomes (8-4)/(-2-0) = 4/-2 = -2 so for Set #1 slope is -2 now you try for Set #2

OpenStudy (anonymous):

12/4

OpenStudy (anonymous):

help me

OpenStudy (binary3i):

yes 12/4 = 3

OpenStudy (anonymous):

so it D???